# Plancraft: an evaluation dataset for planning with LLM agents

### v1

@@ 3.3.1 Dataset Statistics
We define a complexity metric as directly proportional to the number of items used and the number of recipes needed to craft the item. We assign each example trajectory to a complexity bin, from least to most complex (very easy, easy, medium, hard, very hard), and ensure an equal proportion of examples from each bin distribution.
In total, we sample $1145$ training examples, $570$ validation examples, and $580$ test examples. We hold out items such that $79$ items in the validation set are not seen in the training set, and similarly $128$ items in the test set do not appear in the training set, $63$ of which also do not appear in the validation set. Since these evaluation sets are extensive, we also provide smaller subsets of the validation ($110$ examples) and test set ($117$ examples) which we use in our evaluation.
Figure 3: Example flow of Plancraft. The agent can use the varying set of tools as described in its system prompt. The action descriptions and formats in the system prompt are only present if the action is allowed. The dialogue history is passed as a sequence of messages. Environment messages include the task goal and current observation (text or image). If the agent takes an action that does not impact the environment such as think or search, then the environment does not return an observation but a response based on this action.
@@ 3.3.2 Environment Feedback
Since all models in Plancraft can handle text, we propagate formatting errors as text observations. This feedback informs agents about specific issues, allowing agents to adjust predicted outputs accordingly. For example, if an agent attempts to move an item from one slot to another but specifies the same source and destination slots, the environment will return a message indicating that these slots must be different. Or, if an agent generates a quantity that exceeds defined limits, the system provides an error message explaining that the quantity must fall within acceptable parameters. This feedback mechanism allows agents to correct their actions and ensures smoother interactions with the environment.
We implement an early stopping mechanism to accelerate evaluation – if an agent fails to craft any new item after $s=10$ environment steps, then it is considered stuck. We also stop the episode after a maximum number of steps $n=80$ is reached (recall that all examples can be solved in $<30$ steps).
3.3.3 Impossible Tasks
Since we also wish to evaluate whether agents can predict when a task is unsolvable, we include a portion ($20\%$) of the dataset that requires the agent to craft an item, which is impossible given the inventory. To generate this dataset, during sampling, key materials from the path are deliberately removed from the inventory. This is designed to test an agent’s ability to recognise when a task cannot be completed given the available resources. As mentioned, all previous interactive environments (see Table 1) rely on the assumption that the task given to an agent is solvable. However, real-world scenarios are unlikely to meet this assumption and tasks may often be unsolvable due to missing resources or constraints.
To evaluate the agent’s ability on impossible tasks, we introduce the impossible action, which when emitted by an agent will stop the episode. An impossible task is considered successful if and only when the agent emits impossible. For the impossible set of tasks, we report the F1 score of emitting impossible since it is now also possible for the agent to interrupt a solvable task (see Section 4.1).
@@ 3.4 Expert Planner
To establish a baseline, we handcraft a planner to find an efficient sequence of actions required to complete each task. This planner serves as a standard against which the performance of various agent models can be benchmarked.
We implement the solver as a memoized Depth-First search over all possible crafting recipes given an inventory. If found, we return the shortest path between the initial inventory and the target object that needs to be crafted as a series of smelt and move actions. If the solver does not find a shortest path after 30 seconds, we time out the search. Note that for simplicity our planner does not smelt multiple items in a single action and does not move items directly into the crafting grid from the crafting slot [0] or from a smelt action. As a result, it is possible for shorter paths to exist if an agent is able to group multiple smelting actions into one or anticipate the location of where an item needs to be placed after crafting/smelting. The planner allows us to evaluate the length of a generated plan against a handcrafted standard. We also use the solver to set hyper-parameters, such as the maximum number of steps an agent can take to complete a task.
@@ 3.5 Integrating External Knowledge
We implement a search action that allows an LLM to search for a given item in the knowledge base. Since we wish to evaluate the limits of RAG, we design the search operator to return a gold-label instantiation of a recipe required to craft the particular item. Therefore, our RAG mechanism helps estimate the performance of our system under a perfect retriever and parser.
@@ 4 Method
We evaluate both the open-source Llama 3.1 8B and Llama 3.3 70B models (Dubey et al., 2024) as well as gpt-4o-mini222gpt-4o-mini-2024-07-18 (OpenAI, 2024) in our Plancraft environment. All tested models are chat models, and we format the environment observations and feedback as messages from a user role. We show the interactions between the agent and the Plancraft environment in Figure 3. We evaluate each model in the text-only and multi-modal environments, test making available different sets of actions or tools, and measure the impact of zero-shot versus few-shot versus fine-tuning. We implement the ReAct (Yao et al., 2023b) ‘thought’ strategy as a tool, where the agent can emit think as an action within the environment. Any generated text after think is ignored.
Tool use
We evaluate multiple sets of different tools outside of the core environment actions (move and smelt), we test whether including think, an oracle recipe search (RAG) and impossible has an impact on the planning capabilities of agents. If a tool is active, we incorporate a description of it within the system prompt and provide an example of its use in the few-shot examples (see Figure 3).
Fine-Tuning
We fine-tune a Llama 3.1 8B model on oracle planning trajectories from the training set using LoRA (Hu et al., 2021) ($r=64$,$\alpha=32$). These trajectories only consist of Observation, Action pairs directly obtained from our handcrafted planner and only include the smelt and move actions. As a result, we also test whether the fine-tuned LLM can take advantage of new actions that it has not been explicitly trained on.
Image Observations
We evaluate models on both text and image observations. Since the Llama models we evaluate are text-only models, we train a custom pre-trained bounding-box model (Faster R-CNN) to map from images to a list of items, quantities, and positions. We train the bounding-box model separately, sampling random inventories and their accompanying bounding boxes to obtain training data. Using the features of the bounding box, the model is trained to classify not only an object’s class, but also its quantity.
We then map the extracted symbolic representations to the same format as our text-only observations. We use this approach to benchmark how much planning ability is lost if the observation is generated from an imperfect classifier. For gpt-4o-mini, we evaluate using these extracted textual observations from the bounding-box model and also test passing image inputs directly (since gpt-4o-mini supports images).
@@ 4.1 Metrics
For all methods, we report the task success rate – whether or not the goal item has been crafted, the plan length – the number of environment actions the agent has taken (i.e. smelt and move), and the total number of tokens used (both input and output combined). Task success evaluates the percentage of tasks completed successfully by the agent.
Using the expert planner, we can compare an agent’s trajectories compared to those of the expert. We calculate the average Action Efficiency (AE) as the ratio of the number of actions in an agent’s plan $P$ compared to the path of the expert planner $P^{e}$. Note we only consider successful plans in the calculation of this metric:
@@
Lastly, we track the overall compute efficiency using the total number of tokens used as a proxy. [^3] We consider input and output tokens as equivalent. Token usage is heavily impacted by tool use since calling external tools and actions both requires the model to generate additional tokens (to call the tool) but also introduces new information as a result of the external call. However, we note that tools could hide additional compute costs not factored into simple token usage.
@@
M S Llama 8B 0.10 0.00 0.00 0.04 - - - - 13.21 0.86 6.2M
M S Llama 70B 0.37 0.04 0.00 0.16 - - - - 15.94 1.09 8.9M
M S gpt-4o-mini 0.26 0.00 0.00 0.11 - - - - 14.98 0.29 8.2M
M S Llama 8B FT 0.67 0.45 0.12 0.41 - - - - 10.14 0.03 4.1M
M S T Llama 8B 0.15 0.00 0.00 0.06 9.50 - - - 21.80 3.87 11.1M
M S T Llama 70B 0.44 0.15 0.03 0.22 10.57 - - - 21.16 2.50 11.8M
M S T gpt-4o-mini 0.29 0.00 0.00 0.12 13.98 - - - 23.53 2.37 11.6M
M S T Llama 8B FT 0.57 0.42 0.12 0.37 0.06 - - - 10.41 0.07 4.5M
M S T SE Llama 8B 0.57 0.31 0.01 0.30 4.75 3.18 - - 18.41 3.38 9.3M
M S T SE Llama 70B 0.79 0.60 0.23 0.53 6.69 2.28 - - 18.71 4.27 9.8M
M S T SE gpt-4o-mini 0.56 0.21 0.00 0.27 11.16 2.15 - - 21.98 4.91 10.9M
M S T SE Llama 8B FT 0.61 0.39 0.12 0.38 0.00 0.01 - - 10.36 0.11 4.8M
M S T SE I Llama 8B 0.40 0.09 0.01 0.18 2.31 2.36 0.72 0.32 9.77 2.53 4.9M
M S T SE I Llama 70B 0.80 0.65 0.23 0.55 5.20 2.20 0.16 0.64 16.76 4.05 9.6M
M S T SE I gpt-4o-mini 0.56 0.23 0.00 0.28 6.45 1.15 0.48 0.42 14.89 4.94 4.7M
M S T SE I Llama 8B FT 0.58 0.40 0.11 0.36 0.00 0.02 0.00 0.00 10.38 0.09 5.7M
Table 2: Average Success Rates (SR) for different models and sets of actions available: move (M), smelt (S), think (T), search (SE), impossible (I). We also report the average count of special actions used, the average plan length, the Action Efficiency (AE) and the average total number of tokens used. We group very easy and easy difficulty categories together as ‘easy’ and hard and very hard as ‘hard’.
@@
Llama 8B R-CNN 0.00 13.35 3.00 5.3M
Llama 70B R-CNN 0.06 14.47 1.00 6.1M
gpt-4o-mini R-CNN 0.03 14.93 0.23 5.8M
gpt-4o-mini IMG 0.01 15.53 2.60 95.3M
Table 3: Results for different models with the basic sets of actions (M S) and image observations. We use our R-CNN to extract the symbolic representation from images. Since gpt-4o-mini also supports image inputs, we evaluate the model as a unified VLM where observations are passed directly as images (gpt-4o-mini IMG).
@@ 5.1 Act and ReAct baselines
We first restrict the action space to only the environment actions, move (M) and smelt (S), and evaluate the planning capabilities of Llama 3.1 8B, Llama 3.3 70B and gpt-4o-mini in a few-shot setting. We find that gpt-4o-mini and Llama 70B perform similarly, with an overall task success rate of $0.11$ and $0.16$ respectively. Llama 8B performs noticeably worse with a success rate of only $0.04$. In terms of plan efficiency, we find that gpt-4o-mini is closer to the expert planner ($0.29$) than the Llama models ($0.86$ and $1.09$), which indicates that when the model succeeds, it does so efficiently.
We then extend the pool of possible actions to include think, similar to a ReAct strategy (Yao et al., 2023b). Note that we also update the few-shot prompt appropriately to include a description of the action and an example of its use (see Figure 3). We restrict the maximum number of consecutive think actions to three; however, we find that, in practice, the model does not get stuck in a continuous thought loop. Using think, the few-shot models slightly improve in overall task success: Llama 8B $0.06(+0.02)$, Llama 70B $0.22(+0.06)$, gpt-4o-mini $0.12(+0.01)$. However, think also comes with added costs, as the model is now allowed unconstrained generation steps that lead to longer dialogues. We can see the direct effects of this on token usage; for instance, gpt-4o-mini increases its token usage from 8.2M to 11.6M tokens.
5.2 External Knowledge
@@
When we allow search, we enable the models to incorporate the external knowledge contained in the Minecraft Wiki through the RAG oracle retriever. We find that this has substantial benefits for all three few-shot models, boosting Llama 7B to $0.30(+0.24)$, Llama 70B to 0.53(+0.310.53(+0.310.53 ( + 0.31) and gpt-4o-mini to $0.27(+0.15)$ in overall success rate. As with think, search introduces additional steps and therefore we would expect a rise in average tokens used, however, we find that is not the case. Lower token usage can be partially explained by a global decrease in the use of think when search is available. Instead of speculating on how to craft a particular recipe, search allows models to directly retrieve precise information. This reduces the need for additional reasoning steps, resulting in more efficient token usage.
5.3 Predicting Impossible Tasks
@@
When we include impossible as a possibility, we introduce a way for the agent to interrupt an episode. We find that this negatively affects the smaller Llama 8B model ($-0.12$) more than the larger Llama 70B ($+0.02$) and gpt-4o-mini ($+0.01$) in overall success rate. We also record the F1 rate of emitting impossible, and find that the best performing model, the model that can most accurately predict when a task is impossible or not, is the Llama 70B model, with an F1 score of $0.64$. Adding impossible also decreases the usage of tokens, as models can now interrupt an episode when stuck. This reduces the average tokens used by more than half for some models; for example, gpt-4o-mini goes from $10.9$M to $4.7$M in the average tokens used.
5.4 Effects of Fine-tuning
@@
As mentioned in Section 4, we also fine-tune an LLM on expert plans (Llama 8B FT). As in Wang et al. (2022a), we find that smaller models can increase their task success with only about 1k training examples. Llama 8B FT obtains an overall success rate of around $0.41$, which decreases slightly with each additional action we introduce. We find that fine-tuning severely decreases the model’s ability to use new actions. Llama 8B FT almost never uses think, search, and impossible, and thus is also our most efficient model in terms of plan length.
5.5 Image Observations
@@
In Table 3, we use images instead of text inputs. We leverage the pre-trained bounding-box Faster R-CNN model to map images into text descriptions, and confirm a drop in accuracy for all models: Llama 8B ($-0.04$), Llama 70B ($-0.10$), and gpt-4o-mini ($-0.08$). Because gpt-4o-mini supports image inputs, we also evaluate passing images directly to the model instead of the text observations. In this case, the system instructions remain the same, but we modify the few-shot prompt to replace the observations with images. We refer to this setup as gpt-4o-mini IMG and note that it is unable to solve our task ($0.01$) and uses 100 times as many tokens, since every image is converted into a long sequence of tokens.

### v2

@@ 3.3.1 Dataset Statistics
We define a complexity metric as directly proportional to the number of items used and the number of recipes needed to craft the item. We sample trajectories from least to most complex (easy, medium, hard), where hard examples can always be completed in less than $30$ steps (see Appendix A.2 for details). Our complexity correlates with the planner’s search space exploration, with more complex tasks requiring more steps. We stop episodes after a maximum number of steps $n=30$ is reached.
Impossible Tasks. Since we also wish to evaluate whether agents can predict when a task is unsolvable, we include an additional portion ($17\%$) of the dataset that requires the agent to craft an item, which is impossible given the inventory. To generate this dataset, during sampling, key materials from the path are deliberately removed from the inventory. This is designed to test an agent’s ability to recognise when a task cannot be completed given the available resources. As mentioned, all previous interactive environments (see Table 1) rely on the assumption that the task given to an agent is solvable. However, real-world scenarios are unlikely to meet this assumption and tasks may often be unsolvable due to missing resources or constraints.
In total, we sample $1145$ training examples, $570$ validation examples, and $580$ test examples. We hold out items such that $79$ items in the validation set are not seen in the training set, and similarly $128$ items in the test set do not appear in the training set, $63$ of which also do not appear in the validation set. Since these evaluation sets are extensive, we also provide smaller subsets of the validation ($110$ examples) and test set ($117$ examples where $20$ are impossible) which we use in our evaluation.
Figure 3: Example flow of Plancraft. The agent can use the varying set of tools. The dialogue history is passed as a sequence of messages. Environment messages include the task goal and current observation (text or image). If the agent takes an action that does not impact the environment such as think or search, then the environment does not return an observation but a response based on this action.
@@ 3.3.2 Environment Feedback
Since all models in Plancraft can handle text, we propagate formatting errors as text observations. This feedback informs agents about specific issues, allowing agents to adjust accordingly. For example, when an agent tries to move an item between identical source and destination slots, the environment returns a message specifying that different slots must be used. Or, if an agent generates a quantity that exceeds defined limits, the system provides an error message that the quantity must fall within acceptable parameters. This feedback mechanism allows agents to correct their actions and ensures smoother interactions with the environment.
3.3.3 Evaluating Impossible Tasks
To evaluate the agent’s predictions on when a task is impossible, we introduce the impossible action: when emitted by an agent, this stops the episode. An impossible task is considered successful if and only if the agent emits impossible. We report the F1 score of emitting impossible, since it is now possible for the agent to interrupt a solvable task (see Section 4.1).
@@ 3.4 Expert Planner
To evaluate the quality of generated plans (and not just their validity, as measured by success rate), we handcraft a planner to find an efficient sequence of actions required to complete each task. This planner serves as a standard against which the performance of various agent models can be benchmarked.
We implement the solver as a memoized Depth-First search over all possible crafting recipes given an inventory. If found, we return the shortest path between the initial inventory and the target object that needs to be crafted as a series of smelt and move actions. If the solver does not find a shortest path after 30 seconds, we timeout the search. The planner represents the crafting dependency graph as a directed graph where nodes are items and edges represent the crafting relationship between items. This structure allows for efficient planning by focusing only on recipes that are ancestors of the target item in the dependency graph.
Note that, for efficiency, our planner searches over recipes rather than actions. As a result, shorter paths may still exist if an agent can group actions into one. We provide a more detailed discussion of the planner, including pseudocode and its limitations, in Appendix D.
@@ 3.5 Integrating External Knowledge
We implement a search action that allows an LLM to search for a given item in the knowledge base. Since we wish to evaluate the limits of RAG, we design the search operator to return a gold-label instantiation of a recipe required to craft the particular item. The search mechanism uses an exact match to identify the correct recipe for each item, ensuring that the returned information is always accurate. This approach helps estimate the performance of our system under a perfect retriever and parser. Additional details about our RAG implementation are provided in Appendix E.
@@ 4 Method
[…]
Tool use.  We evaluate multiple sets of different tools outside of the core environment actions (move and smelt), we test whether including think, a recipe Oracle Retriever search (RAG) and impossible has an impact on the planning capabilities of agents. If a tool is active, we incorporate a description of it within the system prompt and provide an example of its use in the few-shot examples (see Appendix A.3).
Fine-Tuning.  We fine-tune a Llama 3.1 8B model on $1145$ planning trajectories from the training set using LoRA (Hu et al., 2021) ($r=64$, $\alpha=32$). These trajectories only consist of Observation, Action pairs directly obtained from our handcrafted planner and only include the smelt and move actions. As a result, we also test whether the LoRA fine-tuned LLM can take advantage of new actions that it has not been explicitly trained on.
Image Observations.   We also evaluate models on both text and image inputs. Since Llama models are text-only, we train a Faster R-CNN (Li et al., 2021) to extract item types, quantities, and positions from images (Appendix F). This model is trained separately on randomly sampled inventories. It classifies object types and quantities using bounding box features, achieving $99\%$ slot accuracy and $86\%$ accuracy for correct type and quantity. We then map the extracted symbolic representations to the same format as our text-only observations. The bounding box model benchmarks how much planning ability is lost if an observation is generated from a trained but imperfect visual processing system. For gpt-4o-mini, we evaluate using the predictions from the bounding-box model but also test passing image inputs directly (since gpt-4o-mini supports images). We also evaluate multimodal observations on Qwen 2.5 VL 72B (Bai et al., 2025) and the aforementioned Gemma 3 models.
@@ 4.1 Metrics
For all methods, we report: (a) the task success rate—whether or not the goal item has been crafted; (b) the plan length—the number of environment actions the agent has taken (i.e. smelt and move); and (c) the total number of tokens used (both input and output combined). Task success evaluates the percentage of tasks completed successfully by the agent and excludes the impossible set.
Using the expert planner, we can compare an agent’s trajectories compared to those of the expert. We calculate the average Action Efficiency (AE) as the average difference between the number of actions in an agent’s plan $P$ compared to the length of the path of the expert planner $P^{e}$. Note we only consider successful plans in the calculation of this metric:
@@
Lastly, we track the overall compute efficiency using the total number of tokens used as a proxy. [^3] We consider input and output tokens as equivalent. Token usage is impacted by tool use since calling external tools and actions both requires the model to generate additional tokens (to call the tool) but also introduces new information as a result of the external call. We note that in Plancraft the observation tends to dominate token usage, therefore tool-use decreases the overall token usage since it leads the model to see fewer observations on average. However, we also note that tools could hide additional compute costs not factored into simple token usage analysis.
@@
M S Llama 8B 0.10 0.00 0.00 0.04 - - - - 29.06 4.38 88.9k
M S Llama 70B 0.38 0.12 0.00 0.18 - - - - 25.57 2.38 76.3k
M S gpt-4o-mini 0.16 0.00 0.00 0.07 - - - - 28.12 0.48 82.1k
M S Llama 8B FT 0.63 0.48 0.11 0.40 - - - - 20.29 0.14 56.5k
M S Gemma 12B 0.16 0.03 0.00 0.07 - - - - 28.12 0.85 79.4k
M S Gemma 27B 0.24 0.08 0.01 0.12 - - - - 27.13 1.85 67.5k
M S Qwen3 30B A3B 0.49 0.21 0.04 0.27 - - - - 24.29 3.66 93.6k
M S T Llama 8B 0.14 0.00 0.00 0.06 11.31 - - - 28.80 6.67 68.3k
M S T Llama 70B 0.45 0.30 0.02 0.26 11.79 - - - 24.34 3.32 58.7k
M S T gpt-4o-mini 0.21 0.05 0.01 0.10 15.43 - - - 27.65 3.63 58.4k
M S T Llama 8B FT 0.61 0.48 0.10 0.39 0.07 - - - 20.59 0.26 58.9k
M S T Gemma 12B 0.20 0.06 0.00 0.10 17.53 - - - 27.82 2.85 57.2k
M S T Gemma 27B 0.38 0.20 0.00 0.20 15.56 - - - 25.70 3.56 56.4k
M S T Qwen3 30B A3B 0.34 0.23 0.03 0.20 15.49 - - - 25.61 2.90 73.1k
M S T I Llama 70B 0.45 0.29 0.03 0.26 6.98 - 0.43 0.45 16.81 2.75 40.3k
M S T I Gemma 27B 0.30 0.15 0.00 0.16 2.82 - 0.82 0.34 7.68 2.89 14.0k
M S T I Qwen3 30B A3B 0.29 0.17 0.01 0.16 3.04 - 0.71 0.37 8.72 1.88 25.1k
M S T SE Llama 8B 0.44 0.14 0.02 0.22 6.23 3.73 - - 25.23 4.48 63.0k
M S T SE Llama 70B 0.91 0.83 0.31 0.67 4.80 2.51 - - 16.99 3.95 38.0k
M S T SE gpt-4o-mini 0.46 0.13 0.02 0.23 13.24 1.93 - - 25.46 5.55 55.4k
M S T SE Llama 8B FT 0.62 0.54 0.10 0.41 0.03 0.08 - - 20.31 0.39 60.0k
M S T SE Gemma 12B 0.53 0.30 0.03 0.29 13.09 1.85 - - 23.75 3.36 49.8k
M S T SE Gemma 27B 0.83 0.69 0.08 0.52 8.77 2.75 - - 19.57 4.06 43.6k
M S T SE Qwen3 30B A3B 0.48 0.38 0.07 0.30 11.59 0.79 - - 23.60 2.84 72.9k
M S T SE I Llama 8B 0.28 0.03 0.00 0.12 2.06 1.81 0.85 0.32 8.29 2.90 15.4k
M S T SE I Llama 70B 0.88 0.82 0.30 0.65 4.57 2.42 0.08 0.71 16.30 3.96 36.7k
M S T SE I gpt-4o-mini 0.40 0.13 0.02 0.20 6.90 1.20 0.60 0.39 15.67 5.12 32.6k
M S T SE I Llama 8B FT 0.62 0.55 0.10 0.41 0.01 0.12 - 0.00 20.24 0.38 60.3k
M S T SE I Gemma 12B 0.43 0.28 0.02 0.25 2.91 0.58 0.71 0.36 9.05 2.87 19.1k
M S T SE I Gemma 27B 0.78 0.62 0.07 0.48 2.73 1.22 0.51 0.45 8.90 3.80 17.5k
M S T SE I Qwen3 30B A3B 0.32 0.28 0.03 0.20 1.69 0.15 0.68 0.38 7.88 2.33 25.5k
Table 2: Average Success Rates (SR) for different models and sets of actions available: move (M), smelt (S), think (T), search (SE), impossible (I). We also report the average count of special actions used, the average plan length, the Action Efficiency (AE) and the average total number of tokens used.
@@
Llama 8B R-CNN 0.00 30.00 - 84.5k
Llama 70B R-CNN 0.13 27.03 2.97 79.5k
gpt-4o-mini R-CNN 0.04 28.93 3.31 82.3k
Gemma 3 12B IMG 0.00 29.96 0.00 137.7k
Gemma 3 27B IMG 0.00 29.94 1.00 117.6k
Qwen 2.5 VL 72B IMG 0.01 29.89 0.50 181.1k
gpt-4o-mini IMG 0.02 29.45 2.50 11.6M
Table 3: Results for different models with the basic sets of actions (M S) and image observations. We use our R-CNN to extract the symbolic representation from images. Since gpt-4o-mini and Gemma 3 also supports image inputs, we evaluate these models as a unified VLM where observations are passed directly as images (IMG). We additionally test the Qwen 2.5 VL 72B model on image only observations.
@@ 5 Results
Act and ReAct baselines.   We first restrict the action space to only the environment actions, move (M) and smelt (S), and evaluate the planning capabilities of LLMs in a few-shot setting. We find that Llama 70B outperforms gpt-4o-mini, with overall task success rates of $0.18$ and $0.07$ respectively. Llama 8B performs noticeably worse with a success rate of only $0.04$. Gemma models perform similarly to gpt-4o-mini with success rates of $0.07$ (12B) and $0.12$ (27B), while the Qwen3 30B A3B reasoning model achieves $0.27$. However comparing Qwen3 is the best model when think is not available is not a fair comparaison as Qwen3 is allowed to generate a reasoning trace before answering. In terms of AE, we find that gpt-4o-mini is closer to the expert planner ($0.48$) than the Llama models ($4.38$ and $2.38$), which indicates that when the model succeeds, it does so efficiently.
We then extend the pool of possible actions to include think, similar to a ReAct strategy (Yao et al., 2023b). Note that we also update the few-shot prompt appropriately to include a description of the action and an example of its use (see Appendix A.3). Using think improves overall task success in all few-shot models: Llama 8B $0.06\ (+0.02)$, Llama 70B $0.26\ (+0.08)$, gpt-4o-mini $0.10\ (+0.03)$. The Gemma models also show modest improvements, while Qwen3 30B A3B slightly decreases to $0.20$ despite its reasoning capabilities. Explicitly including a think action in addition to its reasoning trace decreases Qwen3’s overall success rate.
@@
External Knowledge.   When we allow search, we enable the models to incorporate the external knowledge contained in the Minecraft Wiki through the RAG Oracle Retriever. We find that this has substantial benefits for all three few-shot models, boosting Llama 8B to $0.22\ (+0.16)$, Llama 70B to 0.67(+0.410.67\ (+0.410.67 ( + 0.41) and gpt-4o-mini to $0.23\ (+0.13)$ in overall success rate. Gemma 27B achieves $0.52$ with search, demonstrating significant improvement, while Qwen3 30B A3B reaches $0.30$. Instead of speculating on how to craft a particular recipe, search allows models to directly retrieve precise information and where items must be placed to complete a recipe. This reduces the need for additional reasoning steps, resulting in more efficient token usage.
@@
Predicting Impossible Tasks.   Including impossible into the action inventory introduces a way for the agent to interrupt an episode. This negatively affects the overall success rate of all models. Having large impact on smaller Llama 8B model ($-0.10$) and the Qwen3 30B A3B model ($-0.30$), while having a smaller impact on the larger Llama 70B ($-0.02$) and gpt-4o-mini ($-0.03$). We also record the F1 rate of emitting impossible. The Llama 70B model outperforms all other models on predicting impossible tasks (F1 score $0.71$). Adding impossible also decreases token use, since models can now interrupt an episode when stuck. This reduces the average tokens used significantly for some models; for example, Llama 8B’s average token use reduces from $63.0$k to $15.0$k and gpt-4o-mini’s from $55.4$k to $32.6k$. We note that predicting impossible is not trivial without external knowledge, Llama 70B only gets $0.45$ F1 when search is not available.
@@
Effects of Fine-tuning.   As mentioned in Section 4, we also fine-tune an LLM on expert plans (Llama 8B FT). As in Wang et al. (2022a), we find that smaller models can increase their task success with only about 1k training examples. Llama 8B FT obtains an overall success rate of around $0.40$, which is remains steady with each additional action we introduce. We find that fine-tuning severely decreases the model’s ability to use new actions: Llama 8B FT almost never uses think and search actions, and never emits impossible.
@@
Image Observations.   In Table 3, we use images instead of text inputs. We use the pre-trained bounding-box Faster R-CNN model to map images into text descriptions, and confirm a drop in accuracy for all models: Llama 8B ($-0.04$), Llama 70B ($-0.05$), and gpt-4o-mini ($-0.03$). Because gpt-4o-mini supports image inputs, we also evaluate passing images, and no text descriptions, directly to the model. This is gpt-4o-mini IMG: the system instructions remain the same, but the few-shot prompt replaces the observations with images. This model cannot solve our task ($0.02$) and yet uses significantly more tokens, because every image is converted into a long sequence of tokens ($\approx 200$ per $512\times 512$ image patch). Note that token usage in gpt-4o-mini IMG is an overestimation due to vision tokens being priced at the gpt-4o rate (10×10$\times$10 ×). However, the average token usage remains a direct representation of the average cost per episode.
We also run same image only observations for the Gemma 3 12B/27B and Qwen 2.5 VL 72B class of models, and find similarly that accuracy remains 00 for both Gemma 3 sizes, and Qwen 2.5 VL 72B only manages $1\%$ overall success. To solve Plancraft, a VLM needs to discern where and what objects are in an image, how these change over time, and how to map between these and the text locations/types. VLMs like Gemma 3 26B might approximate positions (e.g., ”bottom left”) but fail to map items to exact grid locations. As Plancraft can contain dialogue sequences of up to 30 observations, VLMs need to be able to handle 30 different image inputs. We hypothesize that this is in part due to the pre-training and fine tuning regime of VLMs which are rarely trained on long image sequences nor to ground arbitrary game UIs. Overall, our results also demonstrate that prompting VLMs using few-shot prompting in the same way as with text does not work in Plancraft. Modern VLMs are unable to leverage the few-shot image examples to calibrate their predictions and generalise more broadly to new tasks.